A sample of heights (in inches) of 20 female statistics students is as follows: 68, 60, 68, 71, 68, 69, 66, 65, 66, 61, 69, 69, 72, 67, 66, 67, 71, 60, 66, 70 Find the median height. Round your answer to 1 decimal place, e.g. 0.5 the absolute tolerance is+/-0 Find the first and third quartiles. Round your answers to 1 decimal place, e.g. 0.5. Q3 Find the inter-quartile range. Round your answer to 1 decimal places, e.g. 0.5. IQR- the absolute tolerance is +/-0.1

Where on the box plot would you find the inter-quartile range? The height of the box O The distance between the points O The point at the far left The length of the box
Answer #1

Ans:

Sorted data:

60 60 61 65 66 66 66 66 67 67 68 68 68 69 69 69 70 71 71 72

Median=(67+68)/2=67.5

Q1=66

Q3=69

IQR=Q3-Q1=69-66=3

The length of the box.